Problem 2. Using dimesional analysis and units: a. For two objects with the same shape but different lengths, how does the area scale with the dimension Length? How does the volume scale with Length]? Explain. (Hint: consider the area of a square and the volume of a cube.) b. Lets apply this: Suppose you are a parent whos child needs to take medicine. You are 6 feet tall and your child is 3 feet tall (assume that you and your child have the same proportions). Knowing your own dose, estimate the dose for your child in the following cases. Explain your reasoning! . The medicine is absorbed throughout the bodys volume. Your dose is 400 mg. » The medicine only coats the surface of your childs throat. Your dose is 30 ml
Answer #1

(a)

A=a^2,V=a^3 (Area and Volume of a Cube)

A=[L]^2

V=[L]^3

(b)

(i)

rac {V_c}{V_p}=left (rac {L_c}{L_p} ight )^3

-(3) 4006

V50 mg

So the dose of the child is 50 mg.

(ii)

rac {A_c}{A_p}=left (rac {L_c}{L_p} ight )^2

rac {A_c}{A_p}=left (rac {3}{6} ight )^2

rac {A_c}{30}=left (rac {3}{6} ight )^2

4c 30 ml 7.5 ml 6